7


A firm's total cost function is given     TC = 4 + 8Q − Q2                                

, and total revenue asTR = 50Q − 4Q2,


where Q is the number of units produced and sold (in thousands). How many units should be produced to maximize profit?

 

a.        14

b.        252

c.        7 000

d.        115 904


 

 



1
Expert's answer
2021-06-11T03:00:21-0400

7 A firm's total cost function is given TC = 4 + 8Q − Q2, and total revenue asTR = 50Q − 4Q2,

where Q is the number of units produced and sold (in thousands). How many units should be produced to maximize profit?


Profit = TR - TC = 50Q - 4Q2 - (4 + 8Q - Q2) = 50Q - 4Q2 - 4 - 8Q + Q2 = -4 + 42Q - 3Q2

Profit has maximum than "\frac{d Profit}{d Q} = 42 - 3*2*Q = 0" from here Q = 7 (thousands)

Answer: c